If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are?


Answer by uday1435(57)

Let the velocity of boat in still water = b and velocity of river = r
Downstream velocity = b + r = 27/3 = 24 mph
Upstream velocity = b – r = 60/10 = 10 mph
Adding the above two equations we get 2b = 24 + 10 = 34
So b 34/2 = 17 mph. Plugging this value into the first equation we get
17+ r = 24 ; r = 24 – 17 = 7 mph
Answer is the river flows at 7 miles per hour and the boat goes in still water at the rate of 17 miles per hour



If you still require further clarifications please write to:
udayakumar.t.r@gmail.com